# On the Complexity of Quantum Field Theory

### v1

@@ Acknowledgements
We are immensely grateful to Gal Binyamini for sharing his deep understanding of the relevant mathematical concepts with us. Furthermore, we would like to thank César Agón, Mike Douglas, Gia Dvali, Arno Hoefnagels, Ro Jefferson, Jeroen Monnee, Eran Palti, David Prieto, and Javier Subils for useful discussions and comments. This research is supported, in part, by the Dutch Research Council (NWO) via a Vici grant.
@@ A.1 O-minimality
One of the main goals of tame geometry to find the most general possible collections of functions and sets which satisfy these axioms. At present, there are many known o-minimal structures. Some of the significant examples which are relevant in this work are the following.
@@
    $\mathbb{R}_{\rm Pfaff}$. This structure is generated by Pfaffian functions as defined in the main text. This o-minimal structure is known for its tractable notion of complexity, which we already used implicitly in our first proposal of QFT complexity. If one demands that the domains of the Pfaffian functions are bounded, one obtains the smaller o-minimal structure $\mathbb{R}_{\rm rPfaff}$.
@@ A.2 Sharp o-minimality
At first glance the appearance of an infinite collection of polynomials $P_{\mathcal{F}}$ seems rather abstract and unpractical, but it is implicitly understood in this definition that these polynomials can be calculated explicitly. The purpose of these polynomials is to provide sharp estimates for any finiteness feature associated to definable sets. The definition of o-minimality is such that finiteness always reduces to finiteness of projected one-dimensional connected components, which are controlled by $P_{\mathcal{F}}(\mathcal{D})$.
@@
    $\mathbb{R}_{\rm an,exp}$. This structure is not sharply o-minimal, as noted in [13]. The reason is that this o-minimal structure contains the class of restricted analytic functions, which are too general to admit any polynomial complexity bounds, as required by sharp o-minimality.
@@
    $\mathbb{R}_{\rm rPfaff}$. Using the strong finiteness theorems obeyed by sets constructed from Pfaffian functions, it has been shown that this o-minimal structure admits a complexity filtration satisfying the axioms listed above [15]. For the details of the construction we refer to the original work and [13]. Essentially, the complexity filtration is constructed by demanding that semi-Pfaffian sets satisfy
@@
    where $\mathcal{F}=n$ and $\mathcal{D}$ is the sum of the degrees of the polynomials required in the construction of $f$ and $g$ as Pfaffian functions. It is conjectured that the structure $\mathbb{R}_{\rm Pfaff}$ which contains unrestricted Pfaffian functions is also sharply o-minimal [13].

### v2

@@ Acknowledgements
We are immensely grateful to Gal Binyamini for sharing his deep understanding of the relevant mathematical concepts with us. Furthermore, we would like to thank César Agón, Mike Douglas, Gia Dvali, Arno Hoefnagels, Ro Jefferson, Jeroen Monnee, Eran Palti, David Prieto, and Javier Subils for useful discussions and comments. Furthermore, we thank the referees from JHEP for valuable comments on an earlier version. This research is supported, in part, by the Dutch Research Council (NWO) via a Vici grant.
@@ A.1 O-minimality
One of the main goals of tame geometry is to find the most general possible collections of functions and sets which satisfy these axioms. At present, there are many known o-minimal structures. Some of the significant examples which are relevant in this work are the following.
@@
    $\mathbb{R}_{\rm Pfaff}$. This structure is generated by Pfaffian functions as defined in the main text. This o-minimal structure is known for its tractable notion of complexity, which we already used implicitly in our first proposal of QFT complexity. If one demands that the domains of the Pfaffian functions are compact and analytic on their domain, one obtains the smaller o-minimal structure $\mathbb{R}_{\rm rPfaff}$.
@@ A.2 Sharp o-minimality
At first glance the appearance of an infinite collection of polynomials $P_{\mathcal{F}}$ seems rather abstract and unpractical, but it is implicitly understood in this definition that these polynomials can be calculated explicitly. The purpose of these polynomials is to provide sharp estimates for any finiteness feature associated to definable sets. The definition of o-minimality is such that finiteness always reduces to finiteness of projected one-dimensional connected components, which are controlled by $P_{\mathcal{F}}(\mathcal{D})$. In this way, the axioms guarantee that the finiteness features of tame sets can be quantified by a polynomial in $\mathcal{D}$, depending on $\mathcal{F}$.
@@
    $\mathbb{R}_{\rm an,exp}$. This structure is not sharply o-minimal, as noted in [13]. The reason is that this o-minimal structure contains the class of restricted analytic functions, which are too general to admit any polynomial complexity bounds, as required by sharp o-minimality. An example of a function definable in $\mathbb{R}_{\rm an,exp}$ which does not admit such bounds is the function
    $f(z)=\sum_{k=1}^{\infty}z^{\omega_{k}}$
    defined on the disk of radius $1/2$, where the sequence $\omega_{k}$ is inductively defined by $\omega_{k}=2^{\omega_{k-1}}$ and $\omega_{1}=1$.
@@
    $\mathbb{R}_{\rm rPfaff}$. Using the strong finiteness theorems obeyed by sets constructed from Pfaffian functions, it has been shown that this o-minimal structure admits a complexity filtration satisfying the axioms listed above [15]. For the details of the construction we refer to the original work and [13]. The complexity filtration is such that semi-Pfaffian sets satisfy
@@
    where $\mathcal{F}=n$ and $\mathcal{D}$ is the sum of the degrees of the polynomials required in the construction of $f$ and $g$ as Pfaffian functions. To assign a complexity to all sets in the structure, one has to address some subtleties involving complements [13]. It is conjectured that the structure $\mathbb{R}_{\rm Pfaff}$ which contains unrestricted Pfaffian functions is also sharply o-minimal [13].